OpenStudy (anonymous):

is there anyone actually here tonight or is everyone just "looking around". i need help with a coulomb's law homework question.

OpenStudy (anonymous):

Two charged particles, with charges q1=q and q2=4q, are located at a distance d=2.00cm apart on the x axis. A third charged particle, with charge q3=q, is placed on the x axis such that the magnitude of the force that charge 1 exerts on charge 3 is equal to the force that charge 2 exerts on charge 3. Find the position of charge 3 when q = 2.00nC .

OpenStudy (anonymous):

got as far as F1on3=F2on3 ==== kq^2/x3^2=4kq^2/(x2-x3)^2. i know the k and the q's cancel, but not sure where to go from there.

OpenStudy (calculusfunctions):

Coulomb's law tells us that the electrical force between two charged objects is directly proportional to the product of their charges and inversely proportional to the square of the distance between their centers.

OpenStudy (anonymous):

Thus you have: 1/x3^2=4/(4-x3)^2. one equation one unknown ,so you can solve it and take only positive number which make sense.

OpenStudy (anonymous):

why did the x2 become 4?

OpenStudy (anonymous):

that was wrong I meant d = 0.02m :D I actually don't know what is x2!!